A parallel plate capacitor having a plate area of 0.70 m2 and a plate separation of 1.0 mm is connected to a source with a


voltage of 50 V. Find the capacitance, the charge on the plates, and the energy of the capacitor when there is air


between the plates.

1
Expert's answer
2022-03-03T12:05:26-0500

The capacitance

"C=\frac{\epsilon_0 A}{d}\\
=\frac{8.85*10^{-12}*0.70}{0.001}=6.2*10^{-9}\:\rm F"

The charge on the plates

"q=CV=6.2*10^{-9}*50=3.1*10^{-7}\:\rm C"

The energy of the capacitor

"W=\frac{CV^2}{2}=\frac{6.2*10^{-9}*50^2}{2}=7.8*10^{-6}\:\rm J"